OpenStudy (anonymous):

Purvi and her friends tipped their waiter a total of twelve dollars for dinners that cost nine dollars , eight dollars , eleven dollars and twelve dollars each. Use a percent proportion to solve for the percent tip they left the waiter. Show your work.

OpenStudy (anonymous):

this is flvs 7 grade math

OpenStudy (anonymous):

First total the cost of the dinner, which is 9+8+11+12=40 percentage of tip= 12/40 * 100= 30%

OpenStudy (anonymous):

tia is right, i got it right i my quiz

OpenStudy (anonymous):

no its not whats flvs?

OpenStudy (anonymous):

lol jk

OpenStudy (anonymous):

i do flvs :P

OpenStudy (anonymous):

hey boys!! hahaha :)

OpenStudy (anonymous):

i know i just finished that chapter

OpenStudy (anonymous):

Im about to be finished

OpenStudy (anonymous):

goood luck because its the worst chapter so far

OpenStudy (anonymous):

i kr ! wow kidz these days! wat could we do !